# FinWorld: An All-in-One Open-Source Platform for End-to-End Financial AI Research and Deployment

### v1

@@ E.3. LLMs Agents for Financial AI
Before introducing financial AI agents, it is important to first accept the concept of a Tool-Calling Agent. In the early stages of agent development, action-based agents played a central role. Notable examples include SayCan (Ahn et al., 2022) and ReAct (Yao et al., 2023), which are built around action selection: they interpret high-level instructions via a language model and choose among a set of predefined skills or actions based on feasibility and expected utility. In contrast, agents such as Voyager (Wang et al., 2023) represent the action generation paradigm, where executable behaviors are dynamically synthesized, often in open-ended environments, allowing the agent to build a library of composable skills without being constrained by a fixed action space.
@@
These advancements culminate in the design of our general-purpose task agent AgentOrchestra (Zhang et al., 2025), which is a hierarchical multi-agent framework designed to systematically address the key challenges of generalization, multimodal reasoning, scalability, and collaboration in complex task-solving environments. The framework adopts a two-tier architecture. At the top level, a planning agent is responsible for understanding user tasks, decomposing them into manageable sub-tasks, and orchestrating the workflow by dynamically assigning these sub-tasks to specialized lower-level agents. The lower level consists of several specialized agents: a Deep Analyzer for in-depth analysis of input data and extraction of key insights, a Deep Researcher that conducts comprehensive research and automatically generates research reports or knowledge summaries, and a Browser Use agent that automates browser-based web searches and information extraction to assist the research process. Additionally, a General Tool Calling Agent provides a unified interface for invoking various external tools and APIs through function calling, enabling efficient execution of specific tasks and integration with diverse external services. Through this layered design, the system achieves flexible coordination, efficient information processing, and robust tool integration.
@@ F.1. Configuration Layer
The configuration layer of FinWorld is built on top of mmengine (Contributors, 2022), providing a unified and extensible configuration system based on Python dictionaries. This design allows all experimental parameters such as dataset sources, model architectures, training procedures, and evaluation settings to be defined in a centralized, human-readable, and programmatically modifiable format. By leveraging mmengine’s support for hierarchical and modular configuration, users can construct complex experimental setups through inheritance, parameter overriding, and dynamic imports. This approach enhances reproducibility, transparency, and collaboration while separating configuration logic from implementation details.
@@ F.2. Dataset Layer
Downloader Module. This module is composed of two tightly integrated components. The first is the Market Data Downloader, which provides a unified interface for heterogeneous financial data providers such as FMP (Team, 2025b) and Alpaca (Team, 2025a). It supports multiple temporal resolutions including daily and minute level data, and accommodates diverse data modalities such as OHLCV time series, fundamentals, news, and alternative signals. The component performs schema harmonization, symbol normalization, and calendar alignment across exchanges. It manages authentication, pagination, and provider specific rate limits, and includes robust retry logic and fault tolerance. All ingested records are validated against a canonical schema and enriched with provenance metadata, including provider, version, retrieval timestamp, and transformation history, in order to ensure complete reproducibility. The second component is the LLM Reasoning Dataset Downloader, which offers programmatic access to financial reasoning corpora such as FinQA, professional certification materials including ACCA and CFA, and curated financial knowledge bases. It enforces licensing checks, version pinning, and deterministic train, validation, and test splits. For textual resources, it provides optional document cleaning, de duplication, and citation preservation. Together, these components implement a single abstraction for acquisition from application programming interfaces and databases, with caching and incremental updates that support scalable and repeatable data collection for research.
Processor Module. Building on standardized access, this module implements configurable preprocessing pipelines for both numerical and textual data. For market data, it supports factor construction such as alpha158 (Yang et al., 2020), rolling window feature engineering, calendar and timezone alignment, corporate action adjustment, missing value imputation, outlier handling, and normalization at the asset level or the cross sectional level. Leakage prevention is enforced through proper temporal ordering, expanding windows, and walk forward splits. For reasoning datasets, the module performs tokenization, span alignment, answer normalization, and the construction of chain of thought or rationales when available. It also supports weak supervision signals, label smoothing, and schema mapping across heterogeneous datasets. All pipelines are declared through a concise configuration that is version controlled and executable, enabling exact reproduction of preprocessing steps across experiments. Comprehensive logging records parameters, random seeds, and checksums for every intermediate artifact.
@@ F.3. Model Layer
Machine Learning Models. This module provides a declarative specification for classical models under a uniform schema. Supported classes include linear and logistic regression, decision trees, random forests, and gradient-boosting ensembles (e.g., XGBoost and LightGBM). Each model is described by a consistent input/output contract and an explicit task head (regression, binary classification, or multiclass classification). Structural attributes, such as regularization choices, tree growth policy, and ensemble topology, are expressed as first-class parameters, enabling transparent comparison across alternatives and systematic hyperparameter control. The uniform abstraction ensures that training and scoring pipelines, metrics, and model persistence follow the same conventions regardless of the underlying estimator.
Deep Learning Models. Neural architectures are organized into composable components, layers, and models with stable interfaces. Components include data embedding for financial time series (Wu et al., 2022), patch embedding (Dosovitskiy et al., 2020), positional encodings, attention mechanisms, transformer blocks, and common activation functions. Layers assemble these components into an embedding layer, an encoder, and a decoder that can be mixed and matched without changing external contracts. Networks such as Autoformer, VAE, and GPT-style decoders are realized by assembling the layers under one specification that remains independent of training and inference. This separation supports rapid iteration on architecture variants while preserving consistent preprocessing, batching, and evaluation semantics.
@@
LLMs Models. This module centralizes access to proprietary and open-source LLMs behind a unified interface. It supports seamless switching among commercial providers (e.g., GPT-4.1, Claude-4-Sonnet, Gemini-2.5-Pro) and efficient local models (e.g., Qwen2.5, Qwen3), with standardized controls for context length, sampling behavior, and cost/latency tracking. Built-in capabilities include function calling, tool use, and retrieval-augmented generation tailored to financial documents. The module integrates naturally with agents and downstream models, enabling document understanding, information extraction, and instruction following under the same invocation patterns as other FinWorld models. This symmetry reduces integration overhead and promotes consistent monitoring and evaluation across modalities.

### v2

@@ E.3. LLMs Agents for Financial AI
Before introducing financial AI agents, it is important to first accept the concept of a Tool-Calling Agent. In the early stages of agent development, action-based agents played a central role. Notable examples include SayCan (ahn2022can) and ReAct (yao2023react), which are built around action selection: they interpret high-level instructions via a language model and choose among a set of predefined skills or actions based on feasibility and expected utility. In contrast, agents such as Voyager (wang2023voyager) represent the action generation paradigm, where executable behaviors are dynamically synthesized, often in open-ended environments, allowing the agent to build a library of composable skills without being constrained by a fixed action space.
@@
These advancements culminate in the design of our general-purpose task agent AgentOrchestra (zhang2025agentorchestra), which is a hierarchical multi-agent framework designed to systematically address the key challenges of generalization, multimodal reasoning, scalability, and collaboration in complex task-solving environments. The framework adopts a two-tier architecture. At the top level, a planning agent is responsible for understanding user tasks, decomposing them into manageable sub-tasks, and orchestrating the workflow by dynamically assigning these sub-tasks to specialized lower-level agents. The lower level consists of several specialized agents: a Deep Analyzer for in-depth analysis of input data and extraction of key insights, a Deep Researcher that conducts comprehensive research and automatically generates research reports or knowledge summaries, and a Browser Use agent that automates browser-based web searches and information extraction to assist the research process. Additionally, a General Tool Calling Agent provides a unified interface for invoking various external tools and APIs through function calling, enabling efficient execution of specific tasks and integration with diverse external services. Through this layered design, the system achieves flexible coordination, efficient information processing, and robust tool integration.
@@ F.1. Configuration Layer
The configuration layer of FinWorld is built on top of mmengine (mmengine2022), providing a unified and extensible configuration system based on Python dictionaries. This design allows all experimental parameters such as dataset sources, model architectures, training procedures, and evaluation settings to be defined in a centralized, human-readable, and programmatically modifiable format. By leveraging mmengine’s support for hierarchical and modular configuration, users can construct complex experimental setups through inheritance, parameter overriding, and dynamic imports. This approach enhances reproducibility, transparency, and collaboration while separating configuration logic from implementation details.
@@ F.2. Dataset Layer
Downloader Module. This module is composed of two tightly integrated components. The first is the Market Data Downloader, which provides a unified interface for heterogeneous financial data providers such as FMP (fmp2025) and Alpaca (alpaca2025). It supports multiple temporal resolutions including daily and minute level data, and accommodates diverse data modalities such as OHLCV time series, fundamentals, news, and alternative signals. The component performs schema harmonization, symbol normalization, and calendar alignment across exchanges. It manages authentication, pagination, and provider specific rate limits, and includes robust retry logic and fault tolerance. All ingested records are validated against a canonical schema and enriched with provenance metadata, including provider, version, retrieval timestamp, and transformation history, in order to ensure complete reproducibility. The second component is the LLM Reasoning Dataset Downloader, which offers programmatic access to financial reasoning corpora such as FinQA, professional certification materials including ACCA and CFA, and curated financial knowledge bases. It enforces licensing checks, version pinning, and deterministic train, validation, and test splits. For textual resources, it provides optional document cleaning, de duplication, and citation preservation. Together, these components implement a single abstraction for acquisition from application programming interfaces and databases, with caching and incremental updates that support scalable and repeatable data collection for research.
Processor Module. Building on standardized access, this module implements configurable preprocessing pipelines for both numerical and textual data. For market data, it supports factor construction such as alpha158 (yang2020qlib), rolling window feature engineering, calendar and timezone alignment, corporate action adjustment, missing value imputation, outlier handling, and normalization at the asset level or the cross sectional level. Leakage prevention is enforced through proper temporal ordering, expanding windows, and walk forward splits. For reasoning datasets, the module performs tokenization, span alignment, answer normalization, and the construction of chain of thought or rationales when available. It also supports weak supervision signals, label smoothing, and schema mapping across heterogeneous datasets. All pipelines are declared through a concise configuration that is version controlled and executable, enabling exact reproduction of preprocessing steps across experiments. Comprehensive logging records parameters, random seeds, and checksums for every intermediate artifact.
@@ F.3. Model Layer
Machine Learning Models. This module provides a declarative specification for classical models under a uniform schema. Supported classes include linear and logistic regression, decision trees, random forests, and gradient‑boosting ensembles (e.g., XGBoost and LightGBM). Each model is described by a consistent input/output contract and an explicit task head (regression, binary classification, or multiclass classification). Structural attributes, such as regularization choices, tree growth policy, and ensemble topology, are expressed as first‑class parameters, enabling transparent comparison across alternatives and systematic hyperparameter control. The uniform abstraction ensures that training and scoring pipelines, metrics, and model persistence follow the same conventions regardless of the underlying estimator.
Deep Learning Models. Neural architectures are organized into composable components, layers, and models with stable interfaces. Components include data embedding for financial time series (wu2022timesnet), patch embedding (dosovitskiy2020image), positional encodings, attention mechanisms, transformer blocks, and common activation functions. Layers assemble these components into an embedding layer, an encoder, and a decoder that can be mixed and matched without changing external contracts. Networks such as Autoformer, VAE, and GPT‑style decoders are realized by assembling the layers under one specification that remains independent of training and inference. This separation supports rapid iteration on architecture variants while preserving consistent preprocessing, batching, and evaluation semantics.
@@
LLMs Models. This module centralizes access to proprietary and open‑source LLMs behind a unified interface. It supports seamless switching among commercial providers (e.g., GPT‑4.1, Claude‑4‑Sonnet, Gemini‑2.5‑Pro) and efficient local models (e.g., Qwen2.5, Qwen3), with standardized controls for context length, sampling behavior, and cost/latency tracking. Built‑in capabilities include function calling, tool use, and retrieval‑augmented generation tailored to financial documents. The module integrates naturally with agents and downstream models, enabling document understanding, information extraction, and instruction following under the same invocation patterns as other FinWorld models. This symmetry reduces integration overhead and promotes consistent monitoring and evaluation across modalities.